The heat of solution of potassium acetate (KC2H3O2) in water is -15.30 kJ/mol.

Heat of Solution 


The heat of solution of potassium acetate (KC2H3O2) in water is -15.30 kJ/mol. 

If 16.80 g of KC2H3O2(S) is dissolved in 525.0 mL of water that is initially at 22.70 °C, what will be the final temperature in degrees Celcius) of the resulting aqueous solution? 

(Assume no heat exchange with the surroundings.)

_______ ℃
Answer #1

no of moles of KC2H3O2   = W/G.M.Wt

                                            = 16.8/98.15 = 0.17116 moles

H = -15.3KJ/mole

      = -15.3KJ/mole*0.17116moles

       = -2.618748KJ

       = -2618.748J

q   = 2618.748J

mass of water = volume * density

                        = 525*1 = 525g

q   = mcT

2618.748   = 525*4.184*(t-22.7)

t = 240C

The final temperature   = 240C